OpenStudy (anonymous):

Given the linear equation y= -6x-2, find the y- coordinates of the points (-1,),(-2,), and (1,). Show all work please.Plot these points and graph the linear equation if you can.

OpenStudy (anonymous):

substitute the value of x in the equation given y1= -6 (-1)-2 =4 y2= -6(-2) -2 =10 y3 = -6(1)-2 =-8 So you have get 3 coordinate: (-1, 4), (-2,10), and (1, -8)